OpenStudy (anonymous):

Earthquakes release seismic waves that occur in concentric circles from the epicenter of the earthquake. Suppose a seismograph station determines the epicenter of an earthquake is located 9 kilometers from the station. If the epicenter is located at the origin, write the equation for the circular wave that passes through the station. A. x2 + y2 = 81 B. x2 + y2 = 9 C. (x - 9)2 + (y - 9)2 = 0 D. (x + 9)2 + (y + 9)2 = 0

OpenStudy (ash2326):

Equation of a circle with center (a, b) and radius r is \[(x-a)^2+(y-b)^2=r^2\] here (a, b)= (0, 0) and radius =9 km so \[ (x-0)^2+(y-0)^2=9^2\] or \[x^2+y^2=81\]

OpenStudy (anonymous):

Thank You !

OpenStudy (ash2326):

Welcome :)